OpenStudy (anonymous):

(v-8/v-9)-(v+1/v+9)+(v-27/v^2-81)

OpenStudy (anonymous):

-27/v^2 + v-9/v -99

OpenStudy (anonymous):

-27/v^2 + v-(9/v) -99

OpenStudy (anonymous):

-27/v^2 +(v^2-9)/v -99

OpenStudy (anonymous):

there's alot of ways to writing it

OpenStudy (anonymous):

do u want it under one fraction

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

(v^3-99v^2-9v-27)/(v^2)

OpenStudy (anonymous):

i think that's the answer

myininaya (myininaya):

\[\frac{v-8}{v-9} \cdot \frac{v+9}{v+9} -\frac{v+1}{v+9} \frac{v-9}{v-9}+\frac{v-27}{(v-9)(v+9)}\] \[=\frac{v^2-8v+9v-72-(v^2-9v+v-9)+(v-27)}{(v-9)(v+9)}\] \[=\frac{v^2-v^2+9v-8v+9v-v+v-72+9-27}{(v-9)(v+9)}\] \[=\frac{10v-90}{(v-9)(v+9)}\] is what i got but i think hahd and interpreted the problem differently

OpenStudy (anonymous):

yes, yes i did

myininaya (myininaya):

when you write (v-8/v-9) actually means to me that you are doing \[v-\frac{8}{v}-9\] but if you say (v-8)/(v-9) actually means to me and everyone that you are doing \[\frac{v-8}{v-9}\] eventhough you didn't write it this second way i mentioned i just tried to use esp

OpenStudy (anonymous):

yh

myininaya (myininaya):

oh yeah we could go alittle further on what i have

myininaya (myininaya):

\[\frac{10(x-9)}{(v-9)(v+9)}=\frac{10}{v+9}, v \neq 9\]